A body thrown up with a finite speed is caught
back after 4 sec. The speed of the body with which it is thrown up is
 

Grade:11

1 Answers

sree lakshmi
3 years ago
 
total time =4sec
time required to reach max height =4/2=2sec
velocity at max height v =0
initial velocity u=?
acceleration = -g
then v=u+at 
          =u-gt=0
         u=gt 
            =9.8×2=19.6m/s
 or approx. = 20m/s